silver oxide (Ag2O) decomposes at temperature 300 C yielding metallic silver and oxygen gas. A 1.60 g sample of impure silver oxide yields 0.104 g of oxygen gas. what is the percent by mass of silver oxide in this formula

Grade:11

2 Answers

Vikas TU
6 years ago
Let x be the wight of the pure Ag2O in the 1.60 sample.
The reaction is:
 
2Ag2O  =>         4Ag +     O2
x/232 moles                   
moles of O2 would be => x/(232*2)
Weight of O2 => (x/464)*32
Hence, given is the mass of O2 => 0.0689x = 0.104 => x = 1.509 gms
Hence the % by mass of silver oxide is:
=> (1.509/1.60)*100
=> 94.33 %
This is the percentage of pure Ag2O in impure sample.
Chetna
5 years ago
2AgO—>4Ag+O2
Moles of O2 =0.104/32=0.0325 moles
From the eqn
1 mole of O2 produce 2 mole of AgO
Therefore, 0.0325 moles produce 0.0325×2=0.065 moles  of AgO .
molecularmass of AgO =216+16= 232g
Weight  of AgO = 0.065×232= 1.508g
Mass% of AgO in sample = (wt .of AgO)×100/total wt. Of sample= 1.508×100/1.6= 94.25%